cross multiply the proportion
500x=(x-5)(500+5x)
Cross multiply and solve
600(x+20=(600-x)/x
solve 12(9-x)+12(9+x)=3
finally
x^2+24=(x+1)^2-25

Answer by checkley71(8403)

500x=(x-5)(500+5x) or 500x=500x-2500+5x~2-25x or 5x~2-25x-2500=0 or
(x-25)(x+20)=0 or x=25 & x=-20